Teder Corporation stock currently sells for $95 per share. The market requires a 7 percent return on the firm's stock.

Required : If the company maintains a constant 4 percent growth rate in dividends, what was the most recent dividend per share paid on the stock?
Answer #1

Based on constant growth dividend discount model, according to which

Po - Divi -9 Po = Price of Stock Divi = Estimated Dividends for Next Period r = Required Rate of Return 9 = Growth Rate

Divi 95 = 007 -0.04

Divi =2.85

Div0 * (1 + g) = Div1

Div0 * (1 + 4%) = $2.85

Div0 = $2.74